A mass on a spring has an angular oscillation frequency of 2.81 rad/s. The mass has a maximum displacement (when t = 0 s) of 0.232 m. If the spring constant is 29.8 N/m, what is the potential energy stored in the mass-spring system when t = 1.42 s?

I know the answer is .350J but not sure how to get there
Answer #1

PE = stored = 0.5 k x^2

x = a cos (wt)
a = 0.232, w = 2.81
at t=0, x = a = max displacement

x = 0.232 cos (2.81*1.42) = 0.232 cos (3.99 radian)
x = - 0.153 m
PE = 0.5*29.8*0.153*0.153 = 0.348 J \approx 0.350J